The effectiveness of a blood-pressure drug is being investigated. An experimenter finds that, on average, the reduction in systolic blood pressure is 64.7 for a sample of size 996 and standard deviation 9.8.
Estimate how much the drug will lower a typical patient's systolic blood pressure (using a 98% confidence level).
Enter your answer as a tri-linear inequality accurate to one decimal place (because the sample statistics are reported accurate to one decimal place).

Answer by Boreal(15235)

98% half-t-interval is t(0.99,df=995)*s/qrt(n)=2.330*9.8/sqrt(996)
z would work as well
=0.72
interval is (63.98, 65.42)
or (64.0, 65.4)
inequality is 64.0 < mean < 65.4 units mmHg